### v2

@@ Abstract
In previous articles we argued that if both influence at a distance and super-determinism are considered unacceptable to explain the results of loophole-free Bell tests, a third way is available, called “predictive incompleteness”: it tells that the usual quantum state $\psi$ is incomplete, as long as the measurement context has not been specified. This is in agreement with the general quantum framework called CSM (Contexts, Systems and Modalities), which provides a unified approach for considering quantum systems within classical contexts. In this article we further illustrate how predictive incompleteness is working, by looking at other ideas based on either the Kolmogorovian Censorship (KC), or on superdeterminism and supermeasurements. We show that these possibilities are very advantageously replaced by the use of predictive incompleteness.
I Introduction.
In the scientific literature there are still debates to decide whether quantum mechanics (QM) is contextual ContextualReview ; ContextualPG or noncontextual histories , and whether quantum probabilities fit within the usual Kolmogorovian framework LZ1 ; LZ2 or not holik ; svozil ; book . To some extent, these debates can be considered as matters of definitions, so it is useful to spell out the explicit or implicit assumptions that are behind these different views. A milestone in this debate is the Kolmogorovian Censorship (KC), stating that quantum probabilities can be identified with classical, Kolmogorovian probabilities when the measurement context has been specified, or more generally when a classical probability distribution of measurement contexts has been specified. The KC sounds rather obvious on a physical intuitive basis, and technically it has been demonstrated for a countable number of measurement contexts proof ; redei . It also has some relationship with superdeterminism or supermeasurements of Refs. Sabine ; Palmer as it will appear more clearly below.
@@
Looking in more details, there is some controversy on the physical meaning of the KC, as we will explain now. In particular, Szabó et al LZ1 ; LZ2 claim that QM probabilities are Kolmogorovian, not only in a single context, but also in a loophole-free Bell experiment, where a random choice among four different contexts is implemented aa . Then they conclude that Bell’s inequalities are either irrelevant (when considering the four contexts separately), or not violated, when considering the four contexts together: since each one has a 1/4 probability to occur, Bell’s $S$ parameter (see Annex) is reduced from $2\sqrt{2}$ down to 2/$\sqrt{2}/2$, that is below 2 as it would be expected classically.
More precisely, it is uncontroversial that the probabilities predicted by QM in a given context are Kolmogorovian. and the problem arises when trying to “put together” probabilities predicted by QM in different contexts. Said otherwise, in any given context it is possible to build a Kolmogorovian probability distribution, or equivalently a hidden variable theory; however this distribution must be adapted to the chosen context.
In a loophole-free Bell test aa , the remote random choices of measurements (polarizers orientations) are designed to forbid that this matching may be done by a relativistically causal transmission between the source and the detectors. Despite these barriers, the matching might be obtained either by an instantaneous influence between the source and the detectors (explicit nonlocality), or by assuming that it is pre-established before the actual experiment takes place (super-determinism). One has then the following possible options:
@@
    Rather than considering a probability distribution over the contexts, one may calculate the correlation function in each context, which is the very idea of a loophole-free Bell test, and gather them in Bell’s $S$ value. However, one may argue that the four correlations functions correspond to four different incompatible experiments, and thus bringing their results together is counterfactual, because they cannot be measured simultaneously; then BI cannot be demonstrated. This is a standard answer to the problem, but without further explanation it sounds more dogmatic than scientific; see III.B for an attempted improvement.
@@
    Still an alternative way is to admit that the results in the four contexts can be combined, as it would be the case classically since they apply to the same system; then one gets Bell’s inequalities, which are experimentally violated, so one should explain why. We consider that this is a meaningful question, and there are basically three options, spelled out in detail in entropy :
    3a - in the spirit of option 1 above, the way to violate Bell’s inequalities is by admitting that the system’s parameters and the orientations of the polarizers are not independent variables, despite the fact that these orientations are chosen randomly and independently at a large distance. This can be obtained either by admitting superdeterminism (i.e. denying the possibility of independent random choices durt ), or by admitting a non-local influence between the source and the measurements entropy . These two options have recently been shown to be equivalent blasiak , and they are in our opinion equally undesirable - though they are matter of ontological choice, and cannot be proven wrong.
    3b - the third option, known as predictive incompleteness entropy ; jarrett , is to recognize that the quantum state by itself is not enough to specify the measured probability distribution, as long as the context has not be specified. This contradicts predictive completeness (also called outcome independence), which is a required hypothesis for Bell’s theorem, and therefore BI cannot be demonstrated. This conclusion corresponds to the detailed analysis presented in entropy , and also the more general framework presented in completing to “complete” the usual quantum state by specifying the measurement context. These two papers, as well as the arguments above, are consistent within the general quantum framework called CSM (Contexts, Systems and Modalities) CSM1 ; CSM2 .
III Discussion.
@@ III.2 What about non-measurable subsets ?
This gives a major hint about the meaning of Palmer’s approach: if counterfactual definiteness - that is, treating in statistical calculations the results of unperformed measurements on an equal footing as actual results - is rejected, then Bell’s inequalities don’t hold any more (option 2. above); this is well known and has not ‘been lost in more recent discussions of Bell’s Theorem’. Then the work in Sabine ; Palmer may be seen as a possible way to justify a rejection of counterfactual definiteness Hance , but there is a much simpler one : counterfactual definiteness is not compatible with predictive incompleteness, see Annex and entropy . Actually Ref. Sabine gives a good definition of predictive incompleteness: it is “the lack of information about the measurement outcomes in the wave-function, combined with the fact that an observation in one location can tell us something about the measurement outcome in another location”. Then Sabine is trying to explain this by using supermeasurements, but here we get it without using any non-measurable set of hidden variables - as a consequence of the CSM contextual quantization postulates CSM1 ; CSM2 ; Gleason ; Uhlhorn . Again, acknowledging that the quantum state vector is predictively incomplete may appear disturbing, but it makes perfect sense and removes many “paradoxes” in the CSM framework.
@@
Again, the KC in a single context is easily integrated in the CSM framework, but one should not conclude that quantum probabilities are Kolmogorovian in a classical sense: this would be true in classical physics, because there is only one universal context, but this fails in quantum physics, because there is a continuous infinity of different, incompatible contexts. This point of view is implicit in textbook quantum mechanics, and it is made explicit in the CSM framework, by using operator algebra and infinite tensor products. This framework allows a contextual unification of classical and quantum physics, within a unique macroscopic physical world MP1 ; MP2 ; MP3 ; enigma .
@@
As a final remark, a significant challenge in quantum reconstructions is to state clearly what are the hypotheses or postulates, and what are their consequences. It should also been made clear whether one looks for a fully deductive reasoning, or a partially inductive one, that is an “Inference to the Best Explanation” Gleason ; Uhlhorn . In the CSM point of view a fully deductive approach does not fit, since, quoting Landau landau , ‘quantum mechanics (…) contains classical mechanics as a limiting case, yet at the same time it requires this limiting case for its own formulation’. Then what is desired is not a deduction of the Laws of Nature from some postulates in a mathematical sense, but rather a fully consistent construction, including both classical and quantum physics from the beginning, and clearly separating experimentally based evidence from its mathematical description MP1 ; MP2 ; MP3 ; enigma .
Acknowledgments. The author thanks Mathias Van Den Bossche and Olivier Ezratty for relevant comments on this article, and for many interesting discussions.
[…]

### v3

@@ Abstract
We revisit the status of quantum probabilities in light of Kolmogorovian Censorship (KC) and the Contexts, Systems and Modalities (CSM) framework, and we compare KC‑based frameworks with alternatives such as superdeterminism, supermeasurements, and predictive incompleteness. After briefly recalling the technical content of KC and its scope, we show that KC correctly identifies that probabilities are classical within a fixed measurement context but does not by itself remove the conceptual tension that motivates nonlocal or conspiratorial explanations of Bell‑inequality violations. We argue that predictive incompleteness — the view that the quantum state is operationally incomplete until the measurement context is specified — provides a simple, minimal, and explanatory framework that preserves relativistic locality while matching experimental practice. Finally we clarify logical relations among these positions, highlight the assumptions behind them, and justify the move from Kolmogorov’s to Gleason’s framework for quantum probabilities.
I Introduction
In the scientific literature there are still debates to decide whether quantum mechanics (QM) is contextual ContextualReview ; ContextualPG or noncontextual histories , and whether quantum probabilities fit within the usual Kolmogorovian framework LZ1 ; LZ2 or not holik ; svozil ; book ; andrei . For the sake of completeness, Kolmogorov’s axioms are briefly presented in Appendix A. To some extent, these debates can be considered as matters of definitions, so it is useful to spell out the explicit or implicit assumptions that are behind these different views. A milestone in this debate is the Kolmogorovian Censorship (KC) LZ1 ; proof ; redei , stating that quantum probabilities can be identified with classical, Kolmogorovian probabilities when the measurement context has been specified, or more generally when a classical probability distribution of measurement contexts has been given. The KC sounds rather obvious on a physical intuitive basis, and technically it has been demonstrated for a countable number of measurement contexts proof ; redei . It also has some relationship with superdeterminism or supermeasurements of Refs. Sabine ; Palmer as it will appear more clearly below.
@@ I Introduction
Looking in more details, there is some controversy on the physical meaning of the KC, as we will explain now. In particular, Szabó et al LZ1 ; LZ2 claim that QM probabilities are Kolmogorovian, not only in a single context, but also in a loophole-free Bell experiment, where a random choice among four different contexts is implemented aa . Then they conclude that Bell’s inequalities are either irrelevant (when considering the four contexts separately), or not violated, when considering the four contexts together: since each one has a 1/4 probability to occur, Bell’s $S$ parameter (see Appendix B) is reduced from $2\sqrt{2}$ down to $\sqrt{2}/2$, that is below 2 as it would be expected classically.
More precisely, it is uncontroversial that the probabilities predicted by QM in a given context, or commuting subalgebra, are Kolmogorovian; the problem arises when trying to “gather” probabilities predicted by QM in different contexts. Said otherwise, in any given context it is possible to build a Kolmogorovian probability distribution, or equivalently a hidden variable theory; however this distribution is matched to the actual context.
In a loophole-free Bell test aa , the remote random choices of measurements (polarizers orientations) are designed to forbid that this matching may be done by a relativistically causal transmission between the source and the detectors. Despite these barriers, the matching might be obtained either by an instantaneous influence between the source and the detectors (explicit nonlocality), or by assuming that it is pre-established before the actual experiment takes place (super-determinism). One has then the following options:
@@
    Rather than considering a probability distribution over the contexts, one may calculate the correlation function in each context, which is the very idea of a loophole-free Bell test, and gather them in Bell’s $S$ value. However, one may argue that the four correlations functions correspond to four different incompatible experiments, and thus bringing their results together is counterfactual, because they cannot be measured simultaneously; then BI cannot be demonstrated. This is a standard answer to the problem, but it remains debated; see below for further explanations.
@@
    Still an alternative way is to admit that the results in the four contexts can be combined, as it would be the case classically since they apply to the same system; then one gets Bell’s inequalities, which are experimentally violated, so one should explain why. We consider that this is a meaningful question, and there are basically three options, spelled out in detail in entropy (see also Appendix C):
    3a, 3b - in the spirit of option 1 above, the way to violate Bell’s inequalities is by admitting that the system’s parameters and the orientations of the polarizers are not independent variables, despite the fact that these orientations are chosen randomly and independently at a large distance. This can be obtained either by (3a) admitting superdeterminism i.e. denying the possibility of independent random choices durt , or by (3b)  admitting a non-local influence between the source and the measurements entropy . These two options have recently been shown to be equivalent blasiak , and they are in our opinion equally undesirable - though they are matter of ontological choice, and cannot be proven wrong.
    3c - the third option, known as predictive incompleteness entropy ; jarrett , is to recognize that the quantum state by itself is not enough to specify the measured probability distribution, as long as the context has not be specified. This contradicts predictive completeness (also called outcome independence), which is a required hypothesis for Bell’s theorem, and therefore BI cannot be demonstrated. This conclusion corresponds to the detailed analysis presented in entropy , and also the more general framework presented in completing to “complete” the usual quantum state by specifying the measurement context. These two papers, as well as the arguments above, are consistent within the general quantum framework called CSM (Contexts, Systems and Modalities) CSM1 ; CSM2 .
III Discussion: from Kolmogorov’s axioms to Gleason’s theorem.
@@ III.2 What about non-measurable subsets ?
This gives a major hint about the meaning of Palmer’s approach: if counterfactual definiteness - that is, treating in statistical calculations the results of unperformed measurements on an equal footing as actual results - is rejected, then Bell’s inequalities don’t hold any more (option 2. above); this is well known and has not ‘been lost in more recent discussions of Bell’s Theorem’. Then the work in Sabine ; Palmer may be seen as a possible way to justify a rejection of counterfactual definiteness Hance , but there is a much simpler one : counterfactual definiteness is not compatible with predictive incompleteness, see Appendices and entropy . Actually Ref. Sabine gives a good definition of predictive incompleteness: it is “the lack of information about the measurement outcomes in the wave-function, combined with the fact that an observation in one location can tell us something about the measurement outcome in another location”. Then Sabine is trying to explain this by using supermeasurements, but here we get it without using any non-measurable set of hidden variables - as a consequence of the CSM contextual quantization postulates CSM1 ; CSM2 ; Gleason ; Uhlhorn . Again, acknowledging that the quantum state vector is predictively incomplete may appear disturbing, but it makes perfect sense and ensures that many “paradoxes” are removed in the CSM framework.
III.3 From $\sigma$‑algebras to projection lattices: Gleason’s route and the gluing of contexts
A compact way to understand the formal passage from classical to quantum probability is to observe that the classical domain of events — a Boolean $\sigma$‑algebra, see Appendix — is replaced in quantum theory by the lattice $\mathcal{P}(\mathcal{H})$ of orthogonal projections on a Hilbert space $\mathcal{H}$. In the classical Kolmogorov framework additivity is required for disjoint sets; Gleason’s theorem implements the exact analogue of this requirement in the quantum setting by demanding additivity for mutually orthogonal projections.
Concretely, one considers a function $m:\mathcal{P}(\mathcal{H})\to[0,1]$ satisfying
  1. 1.
    $m(I)=1$ (normalization),
  2. 2.
    if $P_{i}P_{j}=0$ for $i\neq j$ then $m\!\bigl(\sum_{i}P_{i}\bigr)=\sum_{i}m(P_{i})$ (orthogonal additivity).
Gleason’s theorem shows that, for $\dim\mathcal{H}\geq 3$, any such $m$ is of the trace form,
$m(P_{i})=\operatorname{Tr}(\rho P_{i}),$
for a unique density operator $\rho$. Thus the formal move of replacing a classical $\sigma$‑algebra by the projection lattice and replacing disjoint additivity by orthogonal additivity — is precisely the route Gleason follows. This perspective clarifies the sense in which contexts are “glued.” A single measurement context corresponds to a maximal commuting family of projections; restricted to that commuting subalgebra the projection lattice is isomorphic to a classical Boolean algebra and the probabilities are Kolmogorovian.
The nontrivial content of Gleason’s hypotheses is to ask for a single assignment $m$ that is consistent across all such commuting subalgebras simultaneously: the value assigned to a projection does not depend on which orthogonal decomposition (which context) it appears in. The essential fact that a given projector appears in a continuous infinity of context, provided that $\dim\mathcal{H}\geq 3$, is called intertwining of contexts by Gleason. In the CSM framework, a modality is defined as the association of a projector $P_{i}$ (a usual quantum state) and a context (a complete set of orthogonal projectors including $P_{i}$). Sharing a projector is then an equivalence relation for modalities; it is called extravalence CSM2 , and corresponds physically to mutual certainty. This justifies that the probability depends only on the projector (the extravalence class) and not on the embedding context.
Two additional remarks are in order. First, the lattice $\mathcal{P}(\mathcal{H})$ is non‑Boolean (non‑distributive), and this structural difference is the source of contextuality: additivity on orthogonal families is a strictly different requirement from full $\sigma$‑additivity on a Boolean algebra. Second, Gleason’s theorem requires $\dim\mathcal{H}\geq 3$; nevertheless, the two‑dimensional case is included, provided that it is embedded into higher dimensions Gleason . It is actually well known that one can build a classical model for a single qubit, but not for two qubits.
Operationally, Kolmogorovian Censorship and Gleason’s hypotheses are therefore complementary: KC explains why probabilities look classical inside a fixed context, while Gleason characterizes when and how those context‑by‑context classical assignments can be consistently extended to a single quantum probability law on the whole projection lattice.
III.4 A philosophical detour through physical realsim.
We have seen that the KC in a single context is easily integrated in the CSM framework, but one should not conclude that quantum probabilities are Kolmogorovian in a classical sense: this would be true in classical physics, because there is only one universal context, but this fails in quantum physics, because there is a continuous infinity of different, incompatible contexts. This point of view is implicit in textbook quantum mechanics, and it is made explicit in the CSM framework, by using operator algebra and infinite tensor products. This framework allows a contextual unification of classical and quantum physics, within a unique macroscopic physical world ContextualPG ; MP1 ; MP3 ; enigma .
@@ III.3 From $\sigma$‑algebras to projection lattices: Gleason’s route and the gluing of contexts
The Kolmogorovian Censorship is a useful technical observation: quantum probabilities are Kolmogorovian when a measurement context is fixed, but this formal fact does not resolve the deeper interpretational choice exposed by Bell‑type experiments. Superdeterminism and nonlocal hidden‑variable accounts restore a single global joint distribution only at the cost of questionable ontological commitments and scientific methodology. By contrast, predictive incompleteness accepts that the quantum state alone does not determine outcome statistics across incompatible contexts and thereby preserves locality and empirical adequacy without invoking conspiratorial correlations. It also leads straightforwardly to Gleason’s theorem and Born’s rule. Framing the debate in these terms clarifies which extra assumptions are being made in different reconstructions and points to concrete empirical and conceptual criteria that future work should address.
@@
As a final remark, a significant challenge in quantum reconstructions is to state clearly what are the hypotheses or postulates, and what are their consequences. It should also been made clear whether one looks for a fully deductive reasoning, or a partially inductive one, that is an “Inference to the Best Explanation” Gleason ; Uhlhorn . In the CSM point of view a fully deductive approach does not fit, since, quoting Landau landau , ‘quantum mechanics (…) contains classical mechanics as a limiting case, yet at the same time it requires this limiting case for its own formulation’. Then what is desired is not a deduction of the Laws of Nature from some postulates in a mathematical sense, but rather a fully consistent construction, including both classical and quantum physics from the beginning, and clearly separating experimentally based evidence from its mathematical description ContextualPG ; MP1 ; MP3 ; enigma .
Acknowledgements.
P.G. thanks Mathias Van Den Bossche, Olivier Ezratty, Franck Laloë, Roger Balian, and Hervé Zwirn for interesting discussions. Copilot has been used to edit this document.
Appendix A Classical Kolmogorov framework.
This Appendix provides a short, self‑contained account of the classical Kolmogorov probability framework. We start with the basic mathematical objects, then state the axioms, and finally give some general remarks.
A.1 Basic set‑algebraic objects.
(1) $\sigma$‑algebra. Let $\Omega$ be a nonempty set. A family of subsets $\mathcal{F}\subset 2^{\Omega}$ is a $\sigma$‑algebra (tribu) if (i) $\Omega\in\mathcal{F}$, (ii) $A\in\mathcal{F}\Rightarrow A^{c}\in\mathcal{F}$, and (iii) $\{A_{n}\}_{n\geq 1}\subset\mathcal{F}\Rightarrow\bigcup_{n\geq 1}A_{n}\in\mathcal{F}$. Closure under countable unions (and hence countable intersections) is the defining extra structure that distinguishes a $\sigma$‑algebra from weaker set algebras.
(2) Borel $\sigma$‑algebra. If $(X,\tau)$ is a topological space, the Borel $\sigma$‑algebra $\mathcal{B}(X)$ is the smallest $\sigma$‑algebra containing all open sets (equivalently all closed sets). It is the canonical measurable structure tied to topology and is the usual domain for continuous or Borel‑measurable observables. One often passes to the completion of $\mathcal{B}(X)$ with respect to a given measure to include all null subsets.
(3) Boolean algebra of sets. A family $\mathcal{A}\subset 2^{\Omega}$ is a Boolean algebra if it is closed under finite unions, finite intersections and complementation (equivalently under symmetric difference and intersection). Every $\sigma$‑algebra is a Boolean algebra, but a Boolean algebra need not be closed under countable operations; conversely, the $\sigma$‑algebra generated by a Boolean algebra is obtained by closing it under countable unions and complements.
Why $\sigma$‑algebras for probability. Kolmogorov’s modern axioms require a domain that supports countable additivity and limit operations (monotone convergence, dominated convergence, product measures, etc.). Finite closure (Boolean algebra) is insufficient for these analytic tools. Thus probability theory is formulated on $\sigma$‑algebras.
A.2 Kolmogorov axioms in a measure‑theoretic form.
In order to state the Kolmogorov axioms, one needs a sample space $\Omega$, which is the set of all possible outcomes or elementary events. The space $\mathcal{F}$ of all events, taken to be subsets of $\Omega$, must be a $\sigma$-algebra on $\Omega$, as explained above. Then the probability measure $P$ assigns to each event $A\in F$ its probability $P(A)$ with the properties
  1. 1.
    Positivity: $P(A)\geq 0$.
  2. 2.
    Normalization: $P(\Omega)=1$.
  3. 3.
    Countable additivity: For any countable sequence of pairwise disjoint events $(A_{n})_{n\geq 1}\subset\mathcal{F}$,
    $P\Bigl(\bigcup_{n\geq 1}A_{n}\Bigr)=\sum_{n\geq 1}P(A_{n}).$
From these axioms follow the usual consequences (finite additivity, continuity from above and below, monotonicity). This is the standard modern formulation used throughout probability theory.
A.3 Relations and practical remarks.
  • Generation: Given any collection $\mathcal{C}\subset 2^{\Omega}$ (for instance a Boolean algebra of ”elementary” events), there is a smallest $\sigma$‑algebra $\sigma(\mathcal{C})$ containing $\mathcal{C}$; measures are then defined on $\sigma(\mathcal{C})$. In practice one often starts from a Boolean algebra of operational events and passes to the generated $\sigma$‑algebra to enable countable additivity.
  • Borel vs arbitrary $\sigma$‑algebra: The Borel $\sigma$‑algebra is natural when topology matters (observables continuous or Borel measurable). The full power set $2^{\Omega}$ is usually too large (non‑measurable pathologies); working with Borel sets or its completion avoids these issues.
  • Completion: Given $(\Omega,\mathcal{F},P)$, the completion $\overline{\mathcal{F}}$ adds all subsets of PP‑null sets; completed spaces are convenient for almost‑sure statements and avoid dependence on particular representatives of null classes.
  • Random variables and measurability: A map $X:(\Omega,\mathcal{F})\to(S,\mathcal{S})$ is a random variable if $X^{-1}(B)\in\mathcal{F}$ for every $B\in\mathcal{S}$. When $S=\mathbb{R}$ and $\mathcal{S}=\mathcal{B}(\mathbb{R})$ this is the usual Borel measurability requirement.
  • Product spaces: Product $\sigma$‑algebras (generated by measurable rectangles) are essential for constructing joint distributions and for independence; countable closure is again crucial here.
Algebraic summary. A Boolean algebra models finite logical combinations of events; a $\sigma$‑algebra is the countably closed extension required for measure theory; the Borel $\sigma$‑algebra is the canonical choice when topology is present; Kolmogorov probabilities are measures on $\sigma$‑algebras satisfying normalization and countable additivity, and most practical constructions use Borel (or completed Borel) $\sigma$‑algebras as the measurable domain.
In quantum mechanics the axioms of positivity, normalization and countable additivity are still true, but the global event space in no more a $\sigma$-algebra. This has important consequences discussed in the main text.
Appendix B Bell test with random switching.
Consider a Bell test with fast random switching of the polarizers orientations aa between $x$ and $x^{\prime}$ on one side, $y$ and $y^{\prime}$ on the other side, where the results $A(x)$, $A(x^{\prime})$, $B(y)$, $B(y^{\prime})$ are denoted $\pm 1$. The four correlation coefficients like $E(x,y)=\overline{A(x)B(y)}$ are calculated from the number of counts $N_{\pm}(x,y)$, with similar expressions for $(x^{\prime},y),(x,y^{\prime}),(x^{\prime},y^{\prime})$, which are all gathered together during the experiment. Then Bell’s inequalities tell that |S|2|S| $\leq$ 2, where $S=E(x,y)+E(x^{\prime},y)+E(x^{\prime},y^{\prime})-E(x,y^{\prime})$, in conflict with QM that predicts $S_{max}=2\sqrt{2}$.
@@ Acknowledgements.
Appendix C Three options for violating Bell’s inequalities.
  • Superdeterminism (option 3a)
    • Core idea: Measurement settings are not independent of hidden variables; pre-existing correlations determine outcomes and choices.
    • Costs / risks: Undermines statistical independence/free-choice assumption; risks unfalsifiability if correlations are unconstrained; experimental physics becomes conceptually fragile.
    • Benefits / implications: Restores locality in principle; can reproduce quantum statistics with appropriate hidden-variable dynamics.
  • Nonlocality (option 3b)
    • Core idea: Outcomes at one wing depend on spacelike-separated settings/outcomes at the other. This dependance may occur only at the level on non-accessible hidden variables (de Broglie-Bohm theory).
    • Costs / risks: Tension with relativistic causality and signal locality, even if “no-signalling” is preserved; conceptual burden of action-at-a-distance.
    • Benefits / implications: Explains Bell inequality violations within hidden-variable models; widely used narrative in quantum foundations, often backed up by de Broglie-Bohm theory.
  • Predictive incompleteness (option 3c)
    • Core idea: Quantum state is operationally incomplete until a measurement context is specified; the probabilities are Kolmogorovian per context, not globally across incompatible observables.
    • Costs / risks: Requires explicit conditioning on context and non-Boolean event structure; forbids a single joint distribution for incompatible settings; challenges attitudes based on classical realism.
    • Benefits / implications: Preserves relativistic locality without superdeterminism; reframes Bell’s inequalities violation as category errors of counterfactually mixing contexts; matches standard quantum formalism and experimental practice.
To conclude this Appendix, we briefly sketch the arguments from entropy to show the compatibility between predictive incompleteness and no‑signalling. Let’s preserve elementary locality, also called parameter independence, that is the relativistic no‑signalling constraint, and reject any superluminal causal influence from a distant setting to a local outcome. Bell’s inequalities can still be violated, by rejecting predictive completeness, also called outcome independence: the assumption that the operational state $\psi$ (possibly augmented by $\lambda$) alone suffices to determine the conditional probabilities of outcomes across incompatible contexts entropy .
Concretely, this asserts that the conditional probabilities relevant for Bell factorization must be written with explicit reference to the measurement context $C$, in the sense that the specification of $C$ is required to fix the distribution; therefore the factorization hypothesis used in Bell’s theorem fails for the state $\psi$ alone. This failure is not a dynamical signalling mechanism: it is a statement about the insufficiency of $\psi$ to provide context‑independent outcome statistics, while parameter independence and no‑signalling remain intact.